- ~7/31: Trial for DRMI with polarization.
- 7/6~: Trial for mass lock
- Long term:
- 7/1-8/9: FPMI/DRFPMI commissioning
- 8/12-10/4: ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
Report of work
- ALS DARM was succeeded last night with 30Hz RMS. Target is ~20Hz.
- SR3 has some noise H2 OSEM. It is being traced.
- Mode matching OMC using independent laser outside. It will be connected control today.